# Codes Correcting Two Bursts of Exactly $b$ Deletions

### v4

@@ V Non-binary Two-deletion Correcting Codes
In this section, it is always assumed that $q>2$. We aim to give a new construction of $q$-ary two-deletion correcting codes. Recall the definition of $u(\mathbf{x})$ for any $\mathbf{x}\in\Sigma_{q}^{n}$. If $\mathbf{x}^{\prime}\in\mathcal{B}_{2}(\mathbf{x})$, then $u(\mathbf{x}^{\prime})\in\mathcal{B}_{2}(u(\mathbf{x}))$. So if $u(\mathbf{x})$ belongs to a binary two-deletion correcting code, we can recover $u(\mathbf{x})$ from $u(\mathbf{x}^{\prime})$. We further assume that $\mathbf{x}$ (and thus $u(\mathbf{x})$) is $d$-regular. Then Lemma IV.3 asserts that we can either find two runs of $u(\mathbf{x})$ and each of them suffers one deletion, or find a substring of $u(\mathbf{x})$ of short length which suffers two deletions. This enables us to approximately determine error positions in $\mathbf{x}$. For the latter case, we need a $q$-ary two-deletion correcting code, which is given in Lemma II.4. For the former case, we need a $q$-ary single-deletion correcting code. We use the code given in [23]
@@
Let $u(\mathbf{x})_{I_{j}}$ ($1\leq j\leq r\left(u(\mathbf{x})\right)$) be all the runs in $u(\mathbf{x})$. Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(u(\mathbf{x})\right)}=n$ and $p_{j-1}<p_{j}$ for all $j$. Since $u(\mathbf{x})$ is $d$-regular, we have $\left|I_{j}\right|\leq d\log n$. Therefore, by Lemma V.1, $\mathsf{DVT}\left(I_{j}\right)$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(d\log n)+\log q}$. Let $Q\geq\max\{n,N_{1}\}$ be the smallest prime. By Lemma II.6, we have $n\leq Q<2n$. Define
@@
Since $\left|J_{1}\right|\leq 2P=6d\log n$, we can view each $\xi_{1}\left(\mathbf{x}_{J_{j}}\right)$ as an integer in $\left[0,N_{2}-1\right]$, where $N_{2}=2^{7\left\lceil\log q\right\rceil\log(6d\log n)+O_{q}(1)}$ (see Lemma II.4). For $s\in\{0,1\}$, let $K_{s}=\left\{1\leq j\leq\left\lceil n/P\right\rceil-1:j\equiv s\pmod{2}\right\}$ and define
@@
Then $\mathcal{C}_{2}$ is a two-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b_{0}$, $b_{1}$, $c_{0}$ and $c_{1}$, such that the redundancy of $\mathcal{C}_{2}$ is at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$:
@@
The proof of Theorem V.1 follows directly from Lemmas V.2 and V.3. The claim for redundancy follows from Lemma IV.4 and Lemma II.6.
@@ VI Conclusion
A more complex problem is to construct codes capable of correcting at most two bursts of deletions, where each burst has a length at most $b$. Our technique fails in this setup since we can not even know the number of bursts or the length of each burst occurring in a received sequence. We leave this problem for future research.
The following Lemma .1 is derived by slightly modifying the proof of [42, Proposition 3.3.7]. Before showing its proof, we list some facts.
Let $\ln(\cdot)$ be the natural logarithm. Then when $\left|x\right|<1$, we have $\ln(1+x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{x^{i}}{i}=x-\frac{x^{2}}{2}+\frac{x^{3}}{3}-\cdots$.
Fact .1
For all $10\leq x<$ 1, we have $\ln(1-x)\leq-x-\frac{x^{2}}{2}$ and $-\ln(1+x)\leq-x+\frac{x^{2}}{2}$. When $0\leq x\leq\frac{1}{2}$, we have $-\ln(1-x)\leq x+\frac{x^{2}}{2}+2x^{3}$.
@@
When $10\leq x<$ 1, we have $\ln(1-x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{(-x)^{i}}{i}=-\sum_{i=1}^{\infty}\frac{x^{i}}{i}\leq-x-\frac{x^{2}}{2}$. Since $-\ln(1+x)=-x+\frac{x^{2}}{2}-\left(\ln(1+x)-x+\frac{x^{2}}{2}\right)$ and $\ln(1+x)-x+\frac{x^{2}}{2}\geq 0$ for all $x\geq 0$, we have $-\ln(1+x)\leq-x+\frac{x^{2}}{2}$. At last, it is easy to see that $-\ln(1-x)=x+\frac{x^{2}}{2}+x^{3}\left(\frac{1}{3}+\frac{x}{4}+\frac{x^{2}}{5}+\cdots\right)\leq x+\frac{x^{2}}{2}+x^{3}(1+x+x^{2}+\cdots)=x+\frac{x^{2}}{2}+\frac{x^{3}}{1-x}$. When $x\leq\frac{1}{2}$, we have $1-x\geq\frac{1}{2}$ and thus $-\ln(1-x)\leq x+\frac{x^{2}}{2}+2x^{3}$. ∎
Lemma .1
@@
$H_{q}\left(1-\frac{1}{q}-\epsilon\right)$ $=-\left(1-\frac{1}{q}-\epsilon\right)\log_{q}\left(\frac{1-1/q-\epsilon}{q-1}\right)-\left(\frac{1}{q}+\epsilon\right)\log_{q}\left(\frac{1}{q}+\epsilon\right)$ (23)
$=-\log_{q}\left(\frac{1}{q}\left(1-\frac{\epsilon q}{q-1}\right)\right)+\left(\frac{1}{q}+\epsilon\right)\log_{q}\left(\frac{1-(\epsilon q)/(q-1)}{1+\epsilon q}\right)$
$=1-\frac{1}{\ln q}\left[\ln\left(1-\frac{\epsilon q}{q-1}\right)-\left(\frac{1}{q}+\epsilon\right)\ln\left(\frac{1-(\epsilon q)/(q-1)}{1+\epsilon q}\right)\right]$
$=1+\frac{1}{\ln q}\left[-\ln\left(1-\frac{\epsilon q}{q-1}\right)+\left(\frac{1}{q}+\epsilon\right)\ln\left(1-\frac{\epsilon q}{q-1}\right)-\left(\frac{1}{q}+\epsilon\right)\ln\left(1+\epsilon q\right)\right].$
Since $\epsilon<\min\left\{\frac{1}{q},\frac{q-1}{2q}\right\}$, we have $0<\frac{\epsilon q}{q-1}<\frac{1}{2}$ and $10<\epsilon q<$ 1. Then it follows from .1 that $-\ln\left(1-\frac{\epsilon q}{q-1}\right)\leq\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}}$, $\ln\left(1-\frac{\epsilon q}{q-1}\right)\leq-\frac{\epsilon q}{q-1}-\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}$ and $-\ln(1+\epsilon q)\leq-\epsilon q+\frac{\epsilon^{2}q^{2}}{2}$. Now by Equation 23, we have
$H_{q}\left(1-\frac{1}{q}-\epsilon\right)$ $\leq 1+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\left(\frac{1}{q}+\epsilon\right)\left(-\frac{\epsilon q}{q-1}\right.\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\;\;\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.-\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}-\epsilon q+\frac{\epsilon^{2}q^{2}}{2}\right)\right]$
$=1+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\left(\frac{1}{q}+\epsilon\right)\left(-\frac{\epsilon q^{2}}{q-1}+\frac{\epsilon^{2}q^{3}(q-2)}{2(q-1)^{2}}\right)\right]$
$=1+\frac{\epsilon^{3}q^{3}(q^{2}-3q+6)}{2(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}-\frac{\epsilon^{2}q^{2}}{q-1}+\frac{\epsilon^{2}q^{2}(q-2)}{2(q-1)^{2}}\right]$

### v5

@@ V Non-binary Two-deletion Correcting Codes
In this section, it is always assumed that $q>2$. Recall that $\mathcal{B}_{t}(\mathbf{x})$ is the set of sequences obtained from $\mathbf{x}$ by $t$ deletions. We aim to give a new construction of $q$-ary two-deletion correcting codes. Recall the definition of $u(\mathbf{x})$ for any $\mathbf{x}\in\Sigma_{q}^{n}$. If $\mathbf{x}^{\prime}\in\mathcal{B}_{2}(\mathbf{x})$, then $u(\mathbf{x}^{\prime})\in\mathcal{B}_{2}(u(\mathbf{x}))$. So if $u(\mathbf{x})$ belongs to a binary two-deletion correcting code, we can recover $u(\mathbf{x})$ from $u(\mathbf{x}^{\prime})$. We further assume that $\mathbf{x}$ (and thus $u(\mathbf{x})$) is $d$-regular. Then Lemma IV.3 asserts that we can either find two runs of $u(\mathbf{x})$ and each of them suffers one deletion, or find a substring of $u(\mathbf{x})$ of short length which suffers two deletions. This enables us to approximately determine error positions in $\mathbf{x}$. For the latter case, we need a $q$-ary two-deletion correcting code, which is given in Lemma II.4. For the former case, we need a $q$-ary single-deletion correcting code. We use the code given in [23].
@@
Let $u(\mathbf{x})_{I_{j}}$ ($1\leq j\leq r\left(u(\mathbf{x})\right)$) be all the runs in $u(\mathbf{x})$. Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(u(\mathbf{x})\right)}=n$ and $p_{j-1}<p_{j}$ for all $j$. Since $u(\mathbf{x})$ is $d$-regular, we have $\left|I_{j}\right|\leq d\log n$. Therefore, by Lemma V.1, $\mathsf{DVT}\left(\mathbf{x}_{I_{j}}\right)$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(d\log n)+\log q}$. Let $Q\geq\max\{n,N_{1}\}$ be the smallest prime. By the following lemma, we have $n\leq Q<2n$.
Lemma V.2 (Bertrand–Chebyshev theorem)
Define
@@
Let $\xi_{1}(\cdot)$ be the function in Lemma II.4, which can help to correct two deletions in $q$-ary sequences. Since $\left|J_{1}\right|\leq 2P=6d\log n$, we can view each $\xi_{1}\left(\mathbf{x}_{J_{j}}\right)$ as an integer in $\left[0,N_{2}-1\right]$, where $N_{2}=2^{7\left\lceil\log q\right\rceil\log(6d\log n)+O_{q}(1)}$ (see Lemma II.4). For $s\in\{0,1\}$, let $K_{s}=\left\{1\leq j\leq\left\lceil n/P\right\rceil-1:j\equiv s\pmod{2}\right\}$ and define
@@
Then $\mathcal{C}_{2}$ is a two-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there are some $a$, $b_{0}$, $b_{1}$, $c_{0}$ and $c_{1}$, such that the redundancy of $\mathcal{C}_{2}$ is at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$:
@@
The proof of Theorem V.1 follows directly from Lemmas V.3 and V.4. The claim for redundancy follows from Lemma IV.4 and Lemma V.2.
@@ VI Conclusion
In this work, it is required that the two bursts have the same length. Allowing the two deletion-bursts to have different sizes would make the error model more general and applicable to a wider range of practical scenarios. Indeed, such a generalization would be more realistic, as burst deletions in real-world applications (e.g., communication systems or storage systems) may not always have equal lengths. The case where the two bursts have different lengths seems more involved. Extending our work to handle bursts of different sizes is a promising direction for future research.
A more complex problem is to construct codes capable of correcting at most two bursts of deletions, where each burst has a length at most $b$. Our technique fails in these setup since we can not even know the number of bursts or the length of each burst occurring in a received sequence. This problem is also left for future research.
Acknowledgement
The authors express their gratitude to the anonymous reviewers for their detailed and constructive comments which are very helpful to the improvement of the presentation of this paper. The authors would also like to thank Prof. Alberto Ravagnani, the associate editor, for his excellent editorial job.
Appendix A A Lemma
The following Lemma A.1 is derived by slightly modifying the proof of [43, Proposition 3.3.7]. Before showing its proof, we list some facts. Let $\ln(\cdot)$ be the natural logarithm. Then when $\left|x\right|<1$, we have $\ln(1+x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{x^{i}}{i}=x-\frac{x^{2}}{2}+\frac{x^{3}}{3}$-$\cdots$.
Fact A.1
For all $0\leq x<1$, we have $\ln(1-x)\leq-x-\frac{x^{2}}{2}$ and $\ln(1+x)\leq-x+\frac{x^{2}}{2}$. When $0\leq x\leq\frac{1}{2}$, we have $\ln(1-x)\leq x+\frac{x^{2}}{2}+2x^{3}$.
@@ Appendix A A Lemma
When $0\leq x<1$, we have $\ln(1-x)=\sum_{i=1}^{\infty}(-1)^{i-1}\frac{(-x)^{i}}{i}=-\sum_{i=1}^{\infty}\frac{x^{i}}{i}\leq-x-\frac{x^{2}}{2}$. Since $\ln(1+x)=-x+\frac{x^{2}}{2}-\left(\ln(1+x)-x+\frac{x^{2}}{2}\right)$ and $\ln(1+x)-x+\frac{x^{2}}{2}\geq 0$ for all $x\geq 0$, we have $\ln(1+x)\leq-x+\frac{x^{2}}{2}$. At last, it is easy to see that $\ln(1-x)=x+\frac{x^{2}}{2}+x^{3}\left(\frac{1}{3}+\frac{x}{4}+\frac{x^{2}}{5}+\cdots\right)\leq x+\frac{x^{2}}{2}+x^{3}(1+x+x^{2}+\cdots)=x+\frac{x^{2}}{2}+\frac{x^{3}}{1-x}$. When $x\leq\frac{1}{2}$, we have $1-x\geq\frac{1}{2}$ and thus $\ln(1-x)\leq x+\frac{x^{2}}{2}+2x^{3}$. ∎
Lemma A.1
@@
$H_{q}\left(1-\frac{1}{q}-\epsilon\right)$ (24)
$=$ $-\left(1-\frac{1}{q}-\epsilon\right)\log_{q}\left(\frac{1-1/q-\epsilon}{q-1}\right)$
$\quad-\left(\frac{1}{q}+\epsilon\right)\log_{q}\left(\frac{1}{q}+\epsilon\right)$
$=$ $-\log_{q}\left(\frac{1}{q}\left(1-\frac{\epsilon q}{q-1}\right)\right)$
$\quad+\left(\frac{1}{q}+\epsilon\right)\log_{q}\left(\frac{1-(\epsilon q)/(q-1)}{1+\epsilon q}\right)$
$=$ $1-\frac{1}{\ln q}\left[\ln\left(1-\frac{\epsilon q}{q-1}\right)\right$.
$\quad\left.-\left(\frac{1}{q}+\epsilon\right)\ln\left(\frac{1-(\epsilon q)/(q-1)}{1+\epsilon q}\right)\right]$
$=$ $1+\frac{1}{\ln q}\left[-\ln\left(1-\frac{\epsilon q}{q-1}\right)+\left(\frac{1}{q}+\epsilon\right)\ln\left(1-\frac{\epsilon q}{q-1}\right)\right$.
$\quad\quad\quad\quad\quad\left.-\left(\frac{1}{q}+\epsilon\right)\ln\left(1+\epsilon q\right)\right].$
Since $\epsilon<\min\left\{\frac{1}{q},\frac{q-1}{2q}\right\}$, we have $0<\frac{\epsilon q}{q-1}<\frac{1}{2}$ and $0<\epsilon q<1$. Then it follows from A.1 that $\ln\left(1-\frac{\epsilon q}{q-1}\right)\leq\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}}$, $\ln\left(1-\frac{\epsilon q}{q-1}\right)\leq-\frac{\epsilon q}{q-1}-\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}$ and $\ln(1+\epsilon q)\leq-\epsilon q+\frac{\epsilon^{2}q^{2}}{2}$. Now by Equation 24, we have
$H_{q}\left(1-\frac{1}{q}-\epsilon\right)$
$\leq$ $1+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}\right$.
$\quad\quad\left.+\left(\frac{1}{q}+\epsilon\right)\left(-\frac{\epsilon q}{q-1}-\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}-\epsilon q+\frac{\epsilon^{2}q^{2}}{2}\right)\right]$
$=$ $1+\frac{2\epsilon^{3}q^{3}}{(q-1)^{3}\ln q}+\frac{1}{\ln q}\left[\frac{\epsilon q}{q-1}+\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}\right$.
$\quad\quad\quad\quad\quad\quad\quad\left.+\left(\frac{1}{q}+\epsilon\right)\left(-\frac{\epsilon q^{2}}{q-1}+\frac{\epsilon^{2}q^{3}(q-2)}{2(q-1)^{2}}\right)\right]$
$=$ $1+\frac{\epsilon^{3}q^{3}(q^{2}-3q+6)}{2(q-1)^{3}\ln q}$
$\quad\quad\quad\quad\quad+\frac{1}{\ln q}\left[\frac{\epsilon^{2}q^{2}}{2(q-1)^{2}}-\frac{\epsilon^{2}q^{2}}{q-1}+\frac{\epsilon^{2}q^{2}(q-2)}{2(q-1)^{2}}\right]$